Solve each system by the substitution method.
3y-x=0
2y=4x-6

Answer by funmath(2933)

Solve each system by the substitution method.
L1 3y-x=0
L2 2y=4x-6
:
Isolate a variable on one side of the equal sign and substitute it into the other equation for that variable.
I'm going with L1.
3y-x=0
3y-x+x=0+x
3y=x
Let x in L2 be 3y and solve for y:
2y=4(3y)-6
2y=12y-6
2y-12y=12y-12y-6
-10y=-6
-10y/-10=-6/-10
y=3/5 or .6
Now substitute that in for y in L1 and solve for x.
3(3/5)-x=0
9/5-x=0
9/5-x+x=0+x
9/5=x
x=9/5 or 1.8
The solution (x,y)=(9/5,3/5) or (1.8,.6)
Happy Calculating!!!!